A cylindrical can is to be constructed to hold a fixed volume of liquid. The cost of the

material used for the top and bottom of the can is 3 cents per square inch, and the cost of the

material used for the curved side is 2 cents per square inch. Use calculus to derive a simple

relationship between the radius and height of the can that is the least costly to construct.


1
Expert's answer
2021-11-29T14:17:48-0500

Let "r=" rafius of the cylinder, "h=" its height.


"V=\pi r^2h=>h=\dfrac{V}{\pi r^2}"

The cost of the material is


"C=0.03(2\pi r^2)+0.02(2\pi r h)"

Substitute


"C=C(r)=0.03(2\pi r^2)+0.02(2\pi r (\dfrac{V}{\pi r^2}))"

"=0.06\pi r^2+\dfrac{0.04V}{r}"

Find the first derivative


"C'(r)=0.12\pi r-\dfrac{0.04V}{r^2}"

Find the critical number(s)


"C'(r)=0=>0.12\pi r-\dfrac{0.04V}{r^2}=0"

"r=\sqrt[3]{\dfrac{V}{3\pi}}"

If "0<r<\sqrt[3]{\dfrac{V}{3\pi}}, C'(r)<0, C(r)" decreases.


If "r>\sqrt[3]{\dfrac{V}{3\pi}}, C'(r)>0, C(r)" increases.

The function "C(r)" has a local minimum at "r=\sqrt[3]{\dfrac{V}{3\pi}} ."

Since the function "C(r)" has the only extremum. the function "C"has the absolute minimum at "r=\sqrt[3]{\dfrac{V}{3\pi}} ."



"h=\dfrac{V}{\pi r^2}=\sqrt[3]{\dfrac{9V}{\pi}}"